When asked, "Do kids regularly carry weapons in your school?" 14,256
students said yes, and 115,336 students said no. With a 95% level of
confidence, find the margin of error (MOE) associated with the sample
proportions. (Give the answer to one decimal place, if needed.)
I have been working on these types problems for two days and cannot get them
to come out right. Sure would appreciate any help you could give in that
area. My problem is with figuring the square root. I see how it is done in the book, but do not know how to enter it into the calculator. Thank you so much.

Answer by stanbon(75887)

When asked, "Do kids regularly carry weapons in your school?" 14,256
students said yes, and 115,336 students said no. With a 95% level of
confidence, find the margin of error (MOE) associated with the sample
proportions. (Give the answer to one decimal place, if needed.)
I have been working on these types problems for two days and cannot get them
to come out right. Sure would appreciate any help you could give in that
area. My problem is with figuring the square root. I see how it is done in the book, but do not know how to enter it into the calculator.
--------------------
Note: There are many calculators online. Use Google to find
a scientific calculator..
------------------------------------
Your Problem:
sample proportion: 14,256/(14256+115,336) = 0.11
standard error: E = 1.96*sqrt(0.11*0.89/129593] = 0.0017
-----------
95% CI: 0.11-0.0017 < p < 0.11+0.0017
95% CI: 0.1083 < p < 0.1117
----------------------------------
Note: The square root on a calculator is often over the
x^2 button. Or you could use something like 10^(.5) = 3.1623..
as an example of how to get the square root.
==================================
Cheers,
Stan H.